OpenStudy (anonymous):

Fun Fun Fun. Find the general solution of the given differential equation.Give the largest interval I over which the general solution is defined. Determine whether there are any transient terms in the general solution. cos(x)y'+(sin(x))y=1

OpenStudy (anonymous):

\[e^{\int\limits{}_{}p(x)dx}=e^{\int\limits{}_{}\tan(x)dx}\]

OpenStudy (anonymous):

=\[e^{-\ln(\cos(x)+c}\]

OpenStudy (anonymous):

=\[=e^{\ln(\cos(x^{-1})+c}\]

OpenStudy (anonymous):

\[=\cos(x^{-1})+c\]

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{d}{dx}[\cos(x^{-1})y]=\int\limits_{}^{}\frac{1}{\cos(x)}*\cos(x^{-1})dx\]

OpenStudy (anonymous):

\[\cos(x^{-1})y=\int\limits{}_{}\frac{\cos(x^{-1}}{\cos(x)}dx\]

OpenStudy (anonymous):

\[e ^{\int\limits \tan(x)} = e ^{-\ln \cos(x)} = e ^{\ln {1 \over \cos(x)}} = \sec(x)\]

OpenStudy (anonymous):

aha not simplified enough alright

OpenStudy (anonymous):

\[\sec(x)y=\int\limits_{}^{}\sec^2(x)dx\]

OpenStudy (anonymous):

\[\sec(x)y=\tan(x)+c\]

OpenStudy (anonymous):

so far soo good?

OpenStudy (anonymous):

yeah the book just multiplies by cos(x) to get y=sin(x)+c(cos(x) book makes no sense at all -.- leaves it in c= and then sometimes y=

jimthompson5910 (jim_thompson5910):

because sec(x) = 1/cos(x), they multiply both sides by cos(x) to isolate y

jimthompson5910 (jim_thompson5910):

but yes, looks good Outkast3r09

OpenStudy (anonymous):

Looks good to me. This one is easy to solve for y and its usually best to have an explicit solution

OpenStudy (anonymous):

yeah i undrstand that but sometimes it will say differential general and still leave it as c= anyways i was away for cases could you expain how this works jim"?

jimthompson5910 (jim_thompson5910):

not sure what you mean

OpenStudy (anonymous):

well in this equation ther eis only one case but sometimes you can get two cases such as case 1: x cannot equal 0 case 2: x = 0

OpenStudy (anonymous):

it has to do when you divide by something just didn't understand it the way he explained i'm sure i'd understand if he had explained it

jimthompson5910 (jim_thompson5910):

hmm well I see no domain issues with y=sin(x)+c*cos(x), so you don't have to worry about cases there...but...you may have to worry about cases in intermediate steps. I'm not too sure.

OpenStudy (anonymous):

trying to find a case problem hold on

jimthompson5910 (jim_thompson5910):

alright, I'm looking up dealing with cases, but not finding much

OpenStudy (anonymous):

yeah the book doesn't explain it as much but my teacher is a Dr in Differential Equations and always checks every case.. most of the time it's just the general but i've seen problems where there was 2 answers

jimthompson5910 (jim_thompson5910):

hmm alright, let me look some more

jimthompson5910 (jim_thompson5910):

You might be on to something when you say y(0) So y(0)=sin(0)+c*cos(0) ---> y(0) = c And for nonzero x, y(x)=sin(x)+c*cos(x) Another case is x = kpi/2 where k is some integer. If this is the case, then y(x) = (-1)^(k+1) Not sure if I thought of every possible case though.